OpenStudy (anonymous):

Please help, will medal and fan 1 1 2 ____ + ______ = ______ b + 1 b – 1 b2 – 1

OpenStudy (anonymous):

b^2 *

OpenStudy (kmeis002):

We need to try and get rid of the denominators. Can we find the common factors that they share. Hint: Can you factor the R.H.S. denominator?

Directrix (directrix):

@AcademicGurusInc Do not give direct answers on OpenStudy. >>>For this problem, there are no solutions. In other words, the solution(s) does not exist. OpenStudy values the Learning process - not the ‘Give you an answer’ process' Don’t post only answers - guide the asker to a solution. http://openstudy.com/code-of-conduct

OpenStudy (anonymous):

There has to be an answer for it \[\frac{ 1 }{ b+1 }+\frac{ 1 }{ b-1 }=\frac{ 2 }{ b ^{2} -1 }\]

Directrix (directrix):

@AutumnRoseT Why do you say that there has to be an answer? Just wondering before I work on this.

OpenStudy (anonymous):

I already have this question open somewhere else and I"m getting help I closed this question I don't know why its still showing up...

Directrix (directrix):

Closed questions show up after they are closed. Closing them moves them to the "Closed Questions" section which is next to the "Open Questions" tab.

Directrix (directrix):

Closed or not, I am still interested in why there has to be a solution.

OpenStudy (anonymous):

because in my book it says 'find the solution to each equation' and I've never had a question that doesn't have an answer. They never taught us that and it seems unlike the program...

OpenStudy (kmeis002):

Well, it does have an answer: No solution